Solve using matlab and show code:

A reversible chemical reaction 2A + B C can be
Answer #1

a)

Substitute all values

K = (Cc+x) / (Ca-2x)^2(Cb-x)

0.016 = (4+x)/ (42-2x)^2(28-x)

Provide an equation

0.016 (42^2 - 2*42x + 4x^2)*(28-x) = 4+x

((0.016)*(1764) - 84*0.016 x + 0.064x^2)(28-x) = 4+x

28*0.016*1764 - 84*0.016*28x +0.064*28x^2 - 0.016*1764x +  84*0.016x^2 - 0.064x^3 = 4+x

790.272 - 37.63x +1.792x^2 - 28.224x + 1.344x^2 - 0.064x^3 - 4 -x = 0

-0.064x^3 +(1.792+1.344)x^2 + x(-37.63-28.224) + (790.272-4) = 0

-0.064x^3 +3.136x^2 - 65.854x +786.272 = 0

multiply all by -

0.064x^3 - 3.136x^2 +65.854x - 786.272 = 0

solve graphically for "0"

x = 28.58 (approx)

2-666.596 3 -615.206 4 -568.936 5-527.402 6 -490.22 7 -457.006 8 -427.376 9 -400.946 10377.332 11 -356.15 12 -337.016 13 -319

b)

Applying bisection rule...

you will find that 27.5 is the actual vlaue of the method of iteration